# Quantum Corrections to $\eta/s$ from JT Gravity

### v2

@@ 5 Comparison to the absorption cross-section
$\sigma_{\text{abs}}=\frac{4\pi r_{0}^{2}}{\omega}\left(|\langle E_{i}+\omega|\mathcal{O}|E_{i}\rangle|^{2}\rho(E_{i}+\omega)-|\langle E_{i}-\omega|\mathcal{O}|E_{i}\rangle|^{2}\rho(E_{i}-\omega)\right)\,$, (5.3)
where $E_{i}$ is the initial energy of black hole, $\mathcal{O}$ corresponds to the operator dual to the massless scalar field and $\rho$ is the density of black hole states. The papers Mertens:2022irh ; Stanford:2017thb ; Mertens:2017mtv ; Kitaev:2018wpr ; Yang:2018gdb ; Iliesiu:2019xuh calculated the density of black hole states and the matrix elements of conformal primaries in comprehensive studies of the Schwarzian theory and its coupling to conformal matter. Substituting these results into Eq. (5.3), the absorption cross-section [^13] It should be straightforward to generalize the black hole results of Emparan:2025sao to a black brane. reads Emparan:2025sao
$\sigma_{abs}=A_{H}\left(\frac{\sinh\left(2\pi\sqrt{\frac{2(E_{i}+\omega)}{E_{b}}}\right)}{\cosh\left(2\pi\sqrt{\frac{2(E_{i}+\omega)}{E_{b}}}\right)-\cosh\left(2\pi\sqrt{\frac{2E_{i}}{E_{b}}}\right)}-\frac{\sinh\left(2\pi\sqrt{\frac{2(E_{i}-\omega)}{E_{b}}}\right)\Theta(E_{i}-\omega)}{\cosh\left(2\pi\sqrt{\frac{2E_{i}}{E_{b}}}\right)-\cosh\left(2\pi\sqrt{\frac{2(E_{i}-\omega)}{E_{b}}}\right)}\right)\,$, (5.4)
where $A_{H}=4\pi r^{2}_{0}$, $T=\frac{\sqrt{2E_{i}E_{b}}}{2\pi}$ and the quantum scale is $E_{b}=\frac{1}{C}$. At very low frequencies with the ratio $E_{i}/E_{b}$ fixed, Eq. (5.4) reduces to
$\sigma_{\text{abs}}(\omega\rightarrow 0)\rightarrow A_{H}\left(\coth\left(2\pi\sqrt{\frac{2E_{i}}{E_{b}}}\right)+\frac{1}{2\pi}\sqrt{\frac{E_{b}}{2E_{i}}}\right)\,.$ (5.5)
To compare to our expression for $\eta$, we consider the two following regimes:
  • In the semiclassical limit $E_{i}\ll E_{b}$, Eq. (5.5) becomes
    $\sigma_{\text{abs}}(\omega\rightarrow 0)\simeq A_{H}\left(1+\frac{1}{2\pi}\sqrt{\frac{E_{b}}{2E_{i}}}+2\exp\left(-4\pi\sqrt{\frac{2E_{i}}{E_{b}}}\right)+\cdots\right).$ (5.6)
  • In the quantum limit $E_{i}\gg E_{b}$, Eq. (5.5) becomes
    $\sigma_{\text{abs}}(\omega\rightarrow 0)\simeq A_{H}\frac{1}{\pi}\sqrt{\frac{E_{b}}{2E_{i}}}.$ (5.7)
Finally, expressing $E_{i}$ and $E_{b}$ in terms of the temperature scale $T$ and the Schwarzian coupling constant $C$, we find
@@
    $\sigma_{\text{abs}}(\omega\rightarrow 0)\simeq A_{H}\left(1+\frac{1}{4\pi^{2}}\frac{1}{CT}+2e^{-4\pi^{2}CT}+\cdots\right).$ (5.8)
    To leading order in $1/CT$, this behavior is consistent with the temperature dependence obtained from the approximate expression (4.35) of $\eta^{\prime}$ in the semiclassical region $T\gg\frac{1}{C}$. It would be interesting to better understand the origin of the difference at subleading order.
@@
    $\sigma_{\text{abs}}(\omega\rightarrow 0)\simeq A_{H}\frac{1}{2\pi^{2}}\frac{1}{CT}.$ (5.9)
    This is also consistent with the temperature dependence obtained from the approximate expression (4.37) of $\eta^{\prime}$ in the quantum region $T\ll\frac{1}{C}$.
@@ 6 Conclusions
Perhaps the most interesting feature we have observed – a minimum for $\eta/s$ – arises in the semiclassical regime, where the effects of strong quantum fluctuations can be safely neglected. In this temperature range we find that $\eta/s$ deviates from the KSS bound, becoming smaller as $T$ is lowered, until it reaches a minimum at a critical temperature $T_{c}$ given in (4.42). The shear viscosity itself in this region is monotonic. The minimum of $\eta/s$ can be explained by examining the ratio of the quantum corrected entropy $s^{\prime}$ to the tree level entropy $s$. Indeed, $s^{\prime}/s$ is not monotonic, and reaches a maximum in the semi-classical regime, precisely near $T_{c}$. Below the critical temperature, $\eta/s$ starts to grow, and continues to do so rather quickly in the quantum phase. We also find that the temperature behavior of $\eta$ agrees with that of the black brane absorption cross section $\sigma$ computed in Emparan:2025sao , as expected on general grounds given the relation (5.2).
One should keep in mind, however, that at extremely low temperatures the analysis is subtle and several issues arise. First of all, as we have already mentioned, standard notions of hydrodynamics are likely to break down in the deep quantum regime (see e.g. Arean:2020eus ) at temperatures close to zero (see Blake:2017ris ; Crossley:2015evo ; Crossley:2015tka for an approach to quantum hydrodynamics). However, Davison:2013bxa found that a diffusion mode still exists even at zero temperature, and that hydrodynamics might be applicable (see also Edalati:2010pn ). In any case, our computations of $\eta$ in the quantum regime $CT\ll 1$ should be revisited, once we have a proper description of hydrodynamics there. It would be valuable to understand to what extent our $\eta/s$ captures any of the key physics in the quantum phase. Also, as stressed in Emparan:2025sao , when working within the naive hydrodynamic regime while requiring significant quantum fluctuations, the geometry itself may become sub-Planckian, which we don’t know how to describe. Despite these challenges, which we did not attempt to address here, we have included the $CT\ll 1$ computation because it may help guide future holographic studies of transport.
@@
A particularly interesting and challenging open question is the fate of $\eta$, and in turn $\eta/s$, exactly at zero temperature, once the effects of quantum fluctuations in the bulk are taken into account within a proper hydrodynamic framework. While we have computed the shear viscosity with quantum corrections in both semiclassical $CT\gg 1$ and quantum $CT\ll 1$ regimes, it is also interesting to understand the transition between them. This can be done, in principle, by computing Eq. (4.11) numerically. Our analysis could also be extended to other transport coefficients of interest. Finally, we wonder if the absorption cross section can be recovered directly from the geometry – appropriately modifying the metric to take into account quantum effects. We leave these questions to future work.
@@ Acknowledgments
We would like to thank Roberto Emparan, Leo Pando Zayas, Alex Buchel, Cong-Yuan Yue and Mohammad Moezzi for many helpful comments and discussions. The work of S.C. was supported in part by the National Science Foundation under Grant No. PHY-2210271. L.L. was supported in part by the National Natural Science Foundation of China Grants No. 12525503 and No. 12447101.
@@ A.1 Quantum Green’s function for $T\gg\frac{1}{C}$
$\langle\mathcal{G}(t)\rangle$
$=\,$ $\frac{e^{S_{0}}}{Z(\beta)}\frac{1}{2\pi(2C)^{2\ell-2}}\frac{(2\pi)^{2\ell-1}}{\Gamma(2\ell)}\,\int d\omega e^{-i\omega t}$
@@
$\approx\,$ $\frac{1}{e^{\frac{2\pi^{2}C}{\beta}}}\frac{1}{2\pi(2C)^{2\ell-2}}\frac{(2\pi)^{2\ell-1}}{\Gamma(2\ell)}\,\int d\omega e^{-i\omega t}e^{\frac{2\pi^{2}C}{\beta}}$
$\times\textrm{exp}\left[C\omega\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right]\times\left[\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]^{1-2\ell}$
@@
$=\,$ $\frac{1}{2\pi(2C)^{2\ell-2}}\frac{(2\pi)^{2\ell-1}}{\Gamma(2\ell)}\,\int d\omega e^{-i\omega t}\textrm{exp}\left[C\omega\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right]\times\left[\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]^{1-2\ell}$
@@
$=\,$ $2C\,\int\frac{d\omega}{2\pi}e^{-i\omega t}\textrm{exp}\left[\frac{\omega}{2}\times 2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right]\times\left[\frac{2\pi}{2C\left[\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]}\right]^{2\ell-1}$
@@
$=\,$ $2C\,\left[\frac{\pi}{2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right){\rm sinh}\left[\frac{\pi}{2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)}t\right]}\right]^{2\ell}\,$, (A.6)
where in the last equality, we have used the integral method from Lam:2018pvp . Since we want to obtain the retarded Green’s function, following Faulkner:2011tm we need to perform a Fourier transformation. We also need to reinstate the constant $B$ and use the same rescaling of Lam:2018pvp . Finally, we obtain
$\langle\mathcal{G}_{R}(\omega,\beta)\rangle=-(2\pi)^{2\ell-1}\left[2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right]^{1-2\ell}\frac{\Gamma\left(\ell-i\frac{\omega}{2\pi}2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right)}{\Gamma\left(1-\ell-i\frac{\omega}{2\pi}2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right)}\,.$ (A.7)

### v3

@@ 5 Comparison to the absorption cross-section
$\sigma_{\text{abs}}(\omega)=\frac{4\pi r_{0}^{2}}{\omega}\left(|\langle E_{i}+\omega|\mathcal{O}|E_{i}\rangle|^{2}\rho(E_{i}+\omega)-|\langle E_{i}-\omega|\mathcal{O}|E_{i}\rangle|^{2}\rho(E_{i}-\omega)\right)\,$, (5.3)
where $E_{i}$ is the initial energy of black hole, $\mathcal{O}$ corresponds to the operator dual to the massless scalar field and $\rho$ is the density of black hole states. The papers Mertens:2022irh ; Stanford:2017thb ; Mertens:2017mtv ; Kitaev:2018wpr ; Yang:2018gdb ; Iliesiu:2019xuh calculated the density of black hole states and the matrix elements of conformal primaries in comprehensive studies of the Schwarzian theory and its coupling to conformal matter. Substituting these results into Eq. (5.3), the absorption cross-section [^13] It should be straightforward to generalize the black hole results of Emparan:2025sao to a black brane. in the microcanonical ensemble reads Emparan:2025sao
$\sigma^{\rm mic}_{\rm abs}(\omega)=A_{H}\left(\frac{\sinh\left(2\pi\sqrt{\frac{2(E_{i}+\omega)}{E_{b}}}\right)}{\cosh\left(2\pi\sqrt{\frac{2(E_{i}+\omega)}{E_{b}}}\right)-\cosh\left(2\pi\sqrt{\frac{2E_{i}}{E_{b}}}\right)}-\frac{\sinh\left(2\pi\sqrt{\frac{2(E_{i}-\omega)}{E_{b}}}\right)\Theta(E_{i}-\omega)}{\cosh\left(2\pi\sqrt{\frac{2E_{i}}{E_{b}}}\right)-\cosh\left(2\pi\sqrt{\frac{2(E_{i}-\omega)}{E_{b}}}\right)}\right)\,$, (5.4)
where $A_{H}=4\pi r^{2}_{0}$ and the quantum scale is $E_{b}=\frac{1}{C}$. At very low frequencies with the ratio $E_{i}/E_{b}$ fixed, Eq. (5.4) reduces to
$\sigma_{\text{abs}}^{\rm mic}(\omega\rightarrow 0)\rightarrow A_{H}\left(\coth\left(2\pi\sqrt{\frac{2E_{i}}{E_{b}}}\right)+\frac{1}{2\pi}\sqrt{\frac{E_{b}}{2E_{i}}}\right)\,.$ (5.5)
We need to consider the following regimes:
  • the semiclassical limit $E_{i}\ll E_{b}$, for which Eq. (5.5) becomes
    $\sigma_{\text{abs}}^{\rm mic}(\omega\rightarrow 0)\simeq A_{H}\left(1+\frac{1}{2\pi}\sqrt{\frac{E_{b}}{2E_{i}}}+2\exp\left(-4\pi\sqrt{\frac{2E_{i}}{E_{b}}}\right)+\cdots\right).$ (5.6)
  • the quantum limit $E_{i}\gg E_{b}$, for which Eq. (5.5) becomes
    $\sigma_{\text{abs}}^{\rm mic}(\omega\rightarrow 0)\simeq A_{H}\frac{1}{\pi}\sqrt{\frac{E_{b}}{2E_{i}}}.$ (5.7)
To compare our shear viscosity results to the absorption cross-section, we should convert the microcanonical absorption cross-section of Emparan:2025sao to the canonical one. This can be done as follows:
@@
    $\sigma_{\text{abs}}^{\rm can}(\omega)$ $=\frac{1}{Z(\beta)}\int_{0}^{\infty}dE\,\rho(E)e^{-\beta E}\sigma_{\text{abs}}^{\rm mic}$
    $\approx A_{H}\left(1+\frac{1}{4\pi^{2}CT}+\cdots\right),$ (5.8)
    where the black hole density of states $\rho(E)$ and the partition function $Z(\beta)$ are given by
    $\rho(E)$ $=\frac{1}{2\pi^{2}E_{b}}e^{S_{0}}\sinh\left(2\pi\sqrt{\frac{2E}{E_{b}}}\right)\approx\frac{e^{S_{0}}}{4\pi^{2}E_{b}}\exp\left(2\pi\sqrt{\frac{2E}{E_{b}}}\right),$ (5.9)
    $Z(\beta)$ $=\int_{0}^{\infty}dE\,\rho(E)e^{-\beta E}\approx\frac{e^{S_{0}}}{E_{b}^{3/2}\sqrt{2\pi}\beta^{3/2}}\exp\left(\frac{2\pi^{2}}{E_{b}\beta}\right),$ (5.10)
    and we have used the relation $E_{b}=C^{-1}$. We note that (5.8) is consistent with the results of Biggs:2025nzs .
    To leading order in $1/CT$, this behavior is consistent with the temperature dependence obtained from the approximate expression (4.36) of $\eta^{\prime}$ in the semiclassical region $T\gg\frac{1}{C}$.
@@
    $\sigma_{\text{abs}}^{\rm can}(\omega)$ $=\frac{1}{Z(\beta)}\int_{0}^{\infty}dE\,\rho(E)e^{-\beta E}\sigma_{\text{abs}}^{\rm mic}$
    $\approx A_{H}\frac{1}{\pi}\sqrt{\frac{2}{\pi CT}}\,$, (5.11)
    where
    $\rho(E)\approx\frac{1}{\pi E_{b}}\left(\sqrt{\frac{2E}{E_{b}}}\right)\,,\quad Z(\beta)\approx\frac{1}{\sqrt{2\pi}}E_{brk}^{-3/2}\beta^{-3/2}\,.$ (5.12)
    Once again, (5.11) matches the results of Biggs:2025nzs . The expression for the cross section is consistent with the temperature dependence obtained from the approximate expression (4.38) for the shear viscosity $\eta^{\prime}$ in the quantum region $T\ll\frac{1}{C}$.
@@ 6 Conclusions
Perhaps the most interesting feature we have observed – a minimum for $\eta/s$ – arises in the semiclassical regime, where the effects of strong quantum fluctuations can be safely neglected. In this temperature range we find that $\eta/s$ deviates from the KSS bound, becoming smaller as $T$ is lowered, until it reaches a minimum at a critical temperature $T_{c}$ given in (4.43). The shear viscosity itself in this region is monotonic. The minimum of $\eta/s$ can be explained by examining the ratio of the quantum corrected entropy $s^{\prime}$ to the tree level entropy $s$. Indeed, $s^{\prime}/s$ is not monotonic, and reaches a maximum in the semi-classical regime, precisely near $T_{c}$. Below the critical temperature, $\eta/s$ starts to grow, and continues to do so rather quickly in the quantum phase. We also find that the temperature behavior of $\eta$ agrees with that of the black brane absorption cross section $\sigma$ computed in Emparan:2025sao ; Biggs:2025nzs , as expected on general grounds given the relation (5.2).
One should keep in mind, however, that at extremely low temperatures the analysis is subtle and several issues and open questions arise:
  • First of all, as we have already mentioned, standard notions of hydrodynamics are likely to break down in the deep quantum regime (see e.g. Arean:2020eus ) at temperatures close to zero (see Blake:2017ris ; Crossley:2015evo ; Crossley:2015tka for an approach to quantum hydrodynamics). However, Davison:2013bxa found that a diffusion mode still exists even at zero temperature, and that hydrodynamics might be applicable (see also Edalati:2010pn ). In any case, our computations of $\eta$ in the quantum regime $CT\ll 1$ should be revisited, once we have a proper description of hydrodynamics there. It would be valuable to understand to what extent our $\eta/s$ captures any of the key physics in the quantum phase.
  • Also, as stressed in Emparan:2025sao , when working within the naive hydrodynamic regime while requiring significant quantum fluctuations, the geometry itself may become sub-Planckian, which we don’t know how to describe. Despite these challenges, which we did not attempt to address here, we have included the $CT\ll 1$ computation because it may help guide future holographic studies of transport.
@@
  • Another interesting point arises when we consider the KMS condition in the deep quantum regime. The Wightman Green’s function we obtained in (A.15) exhibits a modified period, indicating an emergent effective inverse temperature $\beta_{\rm eff}=\pi\sqrt{2C\beta}$. Since the periodicity of the Green’s function usually determines the structure of the KMS condition, one is naively led to reformulate (4.13) in terms of $\beta_{\rm eff}$. However, doing so leads to an expression for the shear viscosity that does not agree with the cross section in the deep quantum regime (see Appendix A.3 for details). It would be interesting to better understand the physical interpretation of this emergent effective temperature.
@@
  • A particularly interesting and challenging open question is the fate of $\eta$, and in turn $\eta/s$, exactly at zero temperature, once the effects of quantum fluctuations in the bulk are taken into account within a proper hydrodynamic framework. While we have computed the shear viscosity with quantum corrections in both semiclassical $CT\gg 1$ and quantum $CT\ll 1$ regimes, it is also interesting to understand the transition between them. This can be done, in principle, by computing Eq. (4.11) numerically. Our analysis could also be extended to other transport coefficients of interest. Finally, we wonder if the absorption cross section can be recovered directly from the geometry – appropriately modifying the metric to take into account quantum effects.
We leave these questions to future work.
@@ Acknowledgments
We would like to thank Roberto Emparan, Leo Pando Zayas, Alex Buchel, Gustavo J. Turiaci, Hong Liu, Elias Kiritsis, Sameer Murthy, Blaise Goutéraux, Cong-Yuan Yue and Mohammad Moezzi for many helpful comments and discussions. The work of S.C. was supported in part by the National Science Foundation under Grant No. PHY-2210271. L.L. was supported in part by the National Natural Science Foundation of China Grants No. 12525503 and No. 12447101.
@@ A.1 Quantum Green’s function for $T\gg\frac{1}{C}$
$\langle\mathcal{G}(t)\rangle$ $=\,\frac{e^{S_{0}}}{Z(\beta)}\frac{1}{2\pi(2C)^{2\ell-2}}\frac{(2\pi)^{2\ell-1}}{\Gamma(2\ell)}\,\int d\omega e^{-i\omega t}$
@@
$\approx\,$ $\frac{1}{e^{\frac{2\pi^{2}C}{\beta}}}\frac{1}{\pi^{3/2}}\left(\frac{\beta}{2C}\right)^{\frac{3}{2}}\frac{1}{2\pi(2C)^{2\ell-2}}\frac{(2\pi)^{2\ell-1}}{\Gamma(2\ell)}\,\int d\omega e^{-i\omega t}e^{\frac{2\pi^{2}C}{\beta}}$
$\times\textrm{exp}\left[C\omega\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right]\times\left[\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]^{1-2\ell}\times\frac{\sqrt{8C\pi^{3}}}{\beta^{3/2}}$
@@
$=\,$ $\frac{1}{2\pi(2C)^{2\ell-2}C}\frac{(2\pi)^{2\ell-1}}{\Gamma(2\ell)}\,\int d\omega e^{-i\omega t}\textrm{exp}\left[C\omega\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right]\left[\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]^{1-2\ell}$
@@
$=\,$ $2\,\int\frac{d\omega}{2\pi}e^{-i\omega t}\textrm{exp}\left[\frac{\omega}{2}\times 2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right]\times\left[\frac{2\pi}{2C\left[\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]}\right]^{2\ell-1}$
@@
$=\,$ $2\,\left[\frac{\pi}{2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right){\rm sinh}\left[\frac{\pi}{2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)}t\right]}\right]^{2\ell}\,$, (A.6)
where in the last equality, we have used the integral method from Lam:2018pvp . We perform a Fourier transformation of (A.6) and obtain the Wightman function
$\langle\mathcal{G}(\omega)\rangle$ $=\frac{1}{\pi}\textrm{exp}\left[\frac{\omega}{2}\times 2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right]\times\left[\frac{2\pi}{2C\left[\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]}\right]^{2\ell-1}$
$=\frac{1}{{\rm sin}\left(\pi\ell+\pi i\frac{\omega}{2\pi}\times 2C\left[\frac{\beta}{2C}+\frac{(1-2\ell)}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]\right)}\textrm{exp}\left[\frac{\omega}{2}\times 2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right]$
$\times\left[\frac{2\pi}{2C\left[\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]}\right]^{2\ell-1}\frac{\Gamma\left(\ell-i\frac{\omega}{2\pi}\times 2C\left[\frac{\beta}{2C}+\frac{(1-2\ell)}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]\right)}{\Gamma\left(1-\ell-i\frac{\omega}{2\pi}\times 2C\left[\frac{\beta}{2C}+\frac{(1-2\ell)}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]\right)}\,$, (A.7)
where in the last equality we have used the identity
$\Gamma\left(1-\ell-i\frac{\omega}{2\pi}\times 2C\left[\frac{\beta}{2C}+\frac{(1-2\ell)}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]\right)\Gamma\left(\ell+i\frac{\omega}{2\pi}\times 2C\left[\frac{\beta}{2C}+\frac{(1-2\ell)}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]\right)$
$=\frac{\pi}{{\rm sin}\left(\pi\ell+\pi i\frac{\omega}{2\pi}\times 2C\left[\frac{\beta}{2C}+\frac{(1-2\ell)}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]\right)}\,.$ (A.8)
To get the imaginary part of the retarded Green’s function $\langle\mathcal{G}_{R}(\omega,\beta)\rangle$, we use the KMS condition
${\rm Im}\;\langle\mathcal{G}_{R}(\omega,\beta)\rangle=\frac{1}{2}\left(1-e^{-\beta\omega}\right)\langle\mathcal{G}(\omega)\rangle$
$=\frac{1}{2}\left(1-e^{-\beta\omega}\right)\frac{1}{\pi}\frac{\pi}{{\rm sin}\left(\pi\ell+\pi i\frac{\omega}{2\pi}\times 2C\left[\frac{\beta}{2C}+\frac{(1-2\ell)}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]\right)}\textrm{exp}\left[\frac{\omega}{2}\times 2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right]$
$\times\left[\frac{2\pi}{2C\left[\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]}\right]^{2\ell-1}\frac{\Gamma\left(\ell-i\frac{\omega}{2\pi}\times 2C\left[\frac{\beta}{2C}+\frac{(1-2\ell)}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]\right)}{\Gamma\left(1-\ell-i\frac{\omega}{2\pi}\times 2C\left[\frac{\beta}{2C}+\frac{(1-2\ell)}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]\right)}\,.$ (A.9)
Since for positive integer $\ell$, the retarded Green’s function should be purely imaginary, one is free to write the total retarded Green’s function as
$\langle\mathcal{G}_{R}(\omega,\beta)\rangle$ $=i\frac{1}{2}\left(1-e^{-\beta\omega}\right)\frac{\textrm{exp}\left[\frac{\omega}{2}\times 2C\left(\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right)\right]}{{\rm sin}\left(\pi\ell+\pi i\frac{\omega}{2\pi}\times 2C\left[\frac{\beta}{2C}+\frac{(1-2\ell)}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]\right)}$
$\times\left[\frac{2\pi}{2C\left[\frac{\beta}{2C}+\frac{1-2\ell}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]}\right]^{2\ell-1}\frac{\Gamma\left(\ell-i\frac{\omega}{2\pi}\times 2C\left[\frac{\beta}{2C}+\frac{(1-2\ell)}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]\right)}{\Gamma\left(1-\ell-i\frac{\omega}{2\pi}\times 2C\left[\frac{\beta}{2C}+\frac{(1-2\ell)}{2\pi^{2}}\left(\frac{\beta}{2C}\right)^{2}\right]\right)}\,.$ (A.10)